# Training LLMs to Better Self-Debug and Explain Code

### v1

@@ Abstract
In the domain of code generation, self-debugging is crucial. It allows LLMs to refine their generated code based on execution feedback. This is particularly important because generating correct solutions in one attempt proves challenging for complex tasks. Prior works on self-debugging mostly focus on prompting methods by providing LLMs with few-shot examples, which work poorly on small open-sourced LLMs. In this work, we propose a training framework that significantly improves self-debugging capability of LLMs. Intuitively, we observe that a chain of explanations on the wrong code followed by code refinement helps LLMs better analyze the wrong code and do refinement. We thus propose an automated pipeline to collect a high-quality dataset for code explanation and refinement by generating a number of explanations and refinement trajectories and filtering via execution verification. We perform supervised fine-tuning (SFT) and further reinforcement learning (RL) on both success and failure trajectories with a novel reward design considering code explanation and refinement quality. SFT improves the pass@1 by up to 15.92% and pass@10 by 9.30% over four benchmarks. RL training brings additional up to 3.54% improvement on pass@1 and 2.55% improvement on pass@10. The trained LLMs show iterative refinement ability, and can keep refining code continuously. Lastly, our human evaluation shows that the LLMs trained with our framework generate more useful code explanations and help developers better understand bugs in source code.
@@ 1 Introduction
Code generation has become a crucial research task to automatically generate source code based on natural language description [1, 2, 3, 4]. Although the recent Large Language Models (LLMs) have shown impressive capability in code generation, generating the correct code for a complex problem in single attempt is still challenging [5, 6, 7, 8, 9, 10, 11, 12]. This is expected because even for human developers, completing a hard programming problem might need multiple rounds of trial-and-error debugging. Self-debugging capability that allows LLMs to retrospect the incorrect code and make changes to resolve the implementation errors is becoming increasingly more important besides their code generation ability.
Existing works [13, 14] investigate off-the-shelf LLMs in the scale of Codex (code-davinci-002[1], GPT-3.5 and GPT-4, and show that these LLMs are able to self-debug the wrong code they generated via prompting methods in a pipeline of code generation and self-refinement as shown in Figure 1. The user first queries the LLM for a solution for the given programming task and the initial solution from the LLM is verified by executing them against the given unit tests. If the solution passes all the unit tests, it is considered correct. Otherwise, the user collects the unit test feedback and forms a new query to ask the LLM for a refinement. Such a process can iterate until the LLM generates a correct solution or reaches the maximum number of iterations. There are different prompt designs when asking for refinement [13]. Compared with directly asking for a refined solution (referred to as “Code Refinement” in the feedback block), asking LLMs to provide an explanation of the wrong solution and then refine it in a chain-of-thought manner (referred to as “Code Explanation and Refinement” in the feedback block) helps it to better understand the unit test feedback and increases the success rate of providing refined solutions (details in Appendix A.1).
@@
In this work, we propose an automated pipeline to collect a high-quality dataset for code explanation and refinement by generating a number of explanation and refinement trajectories and filtering via execution verification. Supervised fine-tuning (SFT) on the high-quality demonstration data boosts LLMs’ wrong code explanation and code refinement ability with a significant margin compared to existing prompting approaches. Furthermore, we propose reinforcement learning (RL) with a novel reward design that considers code explanation semantics and unit test passing rate, enabling LLMs to generate more useful code explanations and correct refinements. To sum up, this work makes the following contributions:
  • We propose a scalable framework consisting of automated data collection, data verification, supervised fine-tuning, and reinforcement learning with new reward designs, to improve LLMs’ self-debugging ability and generate more helpful code explanations and correct refinements.
  • We perform supervised fine-tuning on three LLMs (StarCoder-15B, CodeLlama-7B, and CodeLlama-13B) with the collected code refinement and wrong code explanation data. Our extensive experiments demonstrate that it significantly improves their wrong code explanation and code refinement ability, with the pass@1 improved by up to 15.92% and the pass@10 improved by up to 9.30% on four benchmarks.
  • We train the fine-tuned LLMs using a reinforcement learning algorithm with a novel reward design considering unit test execution and wrong code explanation semantics, and further improve the pass@1 by up to 3.54% and pass@10 by up to 2.55% over SFT on four benchmarks.
  • We conduct comprehensive ablation and analytical experiments as well as a human evaluation to show the generalizability and effectiveness of our approach.
@@ 2 Approach
In this section, we present our proposed automated pipeline, which includes the collection of high-quality code explanation and refinement data, and the training methods. Figure 2 shows the overview of our approach. We first collect a code explanation and refinement dataset by querying from pretrained or instruct models and verifying its responses with execution feedback to filter and obtain high-quality explanation and refinement data (steps 1 and 2 in Figure 2, Section 2.1). Then the high-quality dataset is used for supervised fine-tuning (step 3 in Figure 2, Section 2.2), which significantly improves the model’s performance in explaining the bug and refining the code. Reinforcement learning with execution feedback is used to further guide the model to generate higher quality response and boost the model performance (step 4 in Figure 2, Section 2.3).
Figure 2: Overview of our approach.
@@ 2.1 Data collection and verification
Table 1 shows the number of correct (passing all the unit tests) and wrong (failing any unit test) solutions sampled for each dataset. For each wrong solution, we need an explanation of the wrong code and a correct refinement to build the code explanation and refinement dataset. We prompt pre-trained or instruction-LLMs with the problem description, wrong solution, and execution feedback (either error message or failed test case) to ask for an explanation and refinement. In this paper, we experimented with GPT-3.5-Turbo and CodeLlama-34B. We take GPT-3.5-Turbo as the example LLM in this section, and an example with it is shown in Appendix A.2. We study the generalization of this data collection with different LLMs in Section 4.3.
@@ 4 Result
To evaluate the effectiveness of our training approach, we study the following research questions (RQs) regarding code generation and code refinement capability, iterative refinement ability, approach generalizability, and the quality of the generated code explanations.
@@ 4.1 RQ1: Code generation and refinement capability
We evaluate the models trained with our approach for their code explanation and refinement ability using four benchmarks: MBPP [3], HumanEval [1], MBPP+ [26], and HumanEval+ [26]. We use pass@k [1] and success refinement rate as the evaluation metric. For the generation of the initial solutions, the models sample 100 solutions per task in the benchmarks (temperature set to 0.8), which are run against the provided test cases. For every incorrect solution that fails any test case, we let the models sample one refinement (and one explanation).
@@ 4.1.1 Pass@k
Table 2 shows the pass@k result on four benchmarks separately. In general, fine-tuning LLMs with our collected code explanation and refinement data significantly improves the pass@1 and pass@10 of all three backbones. For both StarCoder-15B and CodeLlama-13B, RL achieves the best pass@1 and pass@10 (bolded) on all four benchmarks. For CodeLlama-7B, RL achieves the best performance under seven settings except that SFT gets the highest pass@1 on the MBPP benchmark.
Table 2: Pass@k of initial and refined solutions on four benchmarks. Each LLM’s best performance on every benchmark is bolded.
@@
For simpler comparison, Table 3 shows the overall pass@k on MBPP and HumanEval, as well as the improvement brought by SFT and RL. The improvement by SFT is compared to prompting, and the improvement by RL is compared to SFT.
On the overall of MBPP and HumanEval, SFT improves StarCoder-15B by a 15.34% higher pass@1 and a 6.58% higher pass@10 after direct refinement. When asking for code explanation and refinement in the chain of thought, the SFT StarCoder-15B obtains a 15.92% and 6.30% improvement on pass@1 and pass@10. RL training improves SFT by another 2.43% pass@1 and 2.55% pass@10 when directly generating refinement, and improves by another 1.87% pass@1 and 1.85% pass@10 when generating code explanation and refinement. Similar improvement by SFT and RL can also be observed on CodeLlama-7B and 13B backbones.
Table 3: Overall pass@k on MBPP & HumanEval and MBPP+ & HumanEval+. ( ) shows the improvement. SFT’s improvement is compared to prompting, and RL’s is compared to SFT.
StarCoder MBPP & HumanEval MBPP+ & HumanEval+
15B SFT RL SFT RL
Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref.
pass@1 54.35 (+15.34) 54.71 (+15.92) 56.78 (+2.43) 56.58 (+1.87) 49.31 (+12.80) 49.64 (+14.13) 50.87 (+1.56) 51.02 (+1.38)
pass@10 76.23 (+6.58) 76.81 (+6.30) 78.78 (+2.55) 78.66 (+1.85) 71.63 (7.27) 72.36 (+9.30) 72.67 (+1.04) 72.87 (+0.51)
CodeLlama MBPP & HumanEval MBPP+ & HumanEval+
7B SFT RL SFT RL
Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref.
pass@1 57.21 (+14.41) 56.75 (+14.14) 56.62 (-0.60) 57.41 (+0.66) 49.59 (+11.15) 49.95 (+11.55) 50.73 (+1.14) 51.74 (+1.79)
pass@10 78.17 (+5.69) 78.98 (+6.18) 79.21 (+1.04) 80.25 (+1.27) 72.42 (+6.27) 71.81 (+4.74) 74.03 (+1.61) 74.42 (+2.60)
CodeLlama MBPP & HumanEval MBPP+ & HumanEval+
13B SFT RL SFT RL
Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref.
pass@1 60.95 (+12.23) 60.64 (+11.68) 61.91 (+0.96) 61.64 (+1.00) 53.86 (+11.14) 53.26 (+10.51) 57.40 (+3.54) 56.60 (+3.34)
pass@10 80.83 (+3.69) 81.29 (+3.72) 81.09 (+0.26) 81.74 (+0.45) 75.21 (+4.04) 74.87 (+3.02) 76.50 (+1.29) 76.00 (+1.13)
On the overall of MBPP+ and HumanEval+ benchmarks, where each problem has 35$\times$ and 80$\times$ more rigorous test cases [26], we observe larger improvements from RL training on CodeLlama backbones. CodeLlama-7B achieves a 1.79% higher pass@1 and a 2.60% higher pass@10 on the refined solutions with code explanation. CodeLlama-13B achieves a 3.54% higher pass@1 and a 1.29% higher pass@10 on the directly generated refinements. This shows that RL training enables LLMs to generate or refine more robust solutions that can pass rigorous test cases. More experiments and examples of case studies can be found in Appendix A.4.1,  A.5.1,  A.5.2,  A.5.3.
@@ 4.1.2 Success refinement rate
Table 4 lists the success refinement rate generated by each backbone using different approaches. The result is averaged over four benchmarks. For StarCoder-15B, prompting is only able to correctly refine 6.41 – 6.90% of the wrong initial solutions. After being fine-tuned with the collected code explanation and refinement data, the supervised fine-tuned StarCoder-15B refines 16.27 – 16.56% of the wrong solutions with a significant improvement of 9.37 – 10.15%. RL training further increases the refinement rate by 1.03 – 1.23% over SFT.
Table 4: Success refinement rate of different approaches over four benchmarks.
@@
Prompt. SFT RL Prompt. SFT RL
StarCoder-15B 6.90 16.27 (+9.37) 17.50 (+1.23) 6.41 16.56 (+10.15) 17.59 (+1.03)
CodeLlama-7B 8.65 18.14 (+9.49) 19.95 (+1.81) 8.10 17.60 (+9.50) 20.84 (+3.24)
CodeLlama-13B 11.64 18.96 (+7.32) 22.58 (+3.62) 11.97 20.06 (+8.09) 23.50 (+3.44)
The improvement on CodeLlama-7B and 13B backbones is consistent with that on StarCoder-15B, where RL training eventually achieves the highest success refinement rate with a considerable boost of 1.81 – 3.62%.
@@ 4.2 RQ2: Iterative refinement ability
LLMs can self-debug iteratively until they find the correct solutions. Figure 4 shows the overall pass@k of CodeLlama-7B on four benchmarks after at most three rounds of refinements. To simplify the figure, we plot the higher pass@k from "Refine" and "Expl. + Refine" in each round of refinement of each LLM. More results of iterative refinement are in Appendix A.4.2.
@@
Both SFT and RL consistently outperform prompting in all three rounds of refinement. Even with three rounds of refinements, the prompting approach cannot achieve the same pass@k as SFT (e.g., 47.53% vs. 56.75% in Figure 4 (a)) after the first round of refinement. These results show that LLMs trained with our pipeline have iterative refinement ability and can achieve increasingly higher pass@k with more iterations.
4.3 RQ3: Generalizability of approach
To show the generalizability of our framework, i.e., especially that the data collection is independent of GPT-3.5-Turbo, we use CodeLlama-34B as an alternative to GPT-3.5-Turbo for data collection. Since CodeLlama-34B is a pre-trained model, we use few-shot examples in the prompt to make it generate the desired wrong code explanation and refinement. The rest remains the same.
@@ 4.3 RQ3: Generalizability of approach
Table 6: Overall pass@k on MBPP & HumanEval and MBPP+ & HumanEval+, trained with CodeLlama-34B’s data.
CodeLlama MBPP & HumanEval MBPP+ & HumanEval+
7B SFT RL SFT RL
Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref.
pass@1 50.01 (+7.21) 53.15 (+10.54) 53.20 (+0.05) 53.64 (+0.49) 45.62 (+7.18) 46.03 (+7.63) 50.76 (+5.14) 51.31 (+5.28)
pass@10 76.88 (+4.40) 77.00 (+4.20) 77.42 (+0.42) 77.33 (+0.33) 69.57 (+3.42) 69.74 (+2.67) 71.65 (+2.08) 71.98 (+2.24)
Table 5 shows the pass@k results of CodeLlama-7B trained with data collected from CodeLlama-34B and Table 6 shows the overall for simpler comparison. SFT obtains smaller (around 1%-3% less compared to training with GPT-3.5-Turbo’s data) but still notable improvements on the overall pass@1 and pass@10. Furthermore, we observe the same finding that RL training improves more of the performance on MBPP+ and HumanEval+ by 5.28% of pass@1 and 2.24% of pass10. This shows the generalizability of our pipeline and also suggests that collecting data from a stronger LLM benefits the training more in our pipeline. More results are shown in Appendix A.4.3.
4.4 RQ4: Quality of generated explanation
We evaluate if the explanation for wrong code is useful for developers to understand their bugs. We randomly sample 50 problems with initial wrong solutions from the MBPP and HumanEval benchmarks and query different LLMs to generate explanations for the wrong code. We score each code explanation on a scale of 1 – 5 considering their correctness and helpfulness, with 1 being totally incorrect or misleading, and 5 being not only correct but also providing a detailed hint of how to fix the wrong code.
We perform scoring using both GPT-4 and human developers as judges. When using GPT-4 as a rater, we follow previous work [27] to prompt it to score each explanation, and the result is shown in Table 7. Both SFT and RL improve the quality of explanation over the prompting approach. Overall, GPT-4 gives higher scores to the LLMs trained with our approach than the prompting approach. Compared with GPT-3.5-Turbo explanations, the gap is significantly reduced after the training.

### v2

# LeDex: Training LLMs to Better Self-Debug and Explain Code

@@ Abstract
In the domain of code generation, self-debugging is crucial. It allows LLMs to refine their generated code based on execution feedback. This is particularly important because generating correct solutions in one attempt proves challenging for complex tasks. Prior works on self-debugging mostly focus on prompting methods by providing LLMs with few-shot examples, which work poorly on small open-sourced LLMs. In this work, we propose LeDex, a training framework that significantly improves the self-debugging capability of LLMs. Intuitively, we observe that a chain of explanations on the wrong code followed by code refinement helps LLMs better analyze the wrong code and do refinement. We thus propose an automated pipeline to collect a high-quality dataset for code explanation and refinement by generating a number of explanations and refinement trajectories from the LLM itself or a larger teacher model and filtering via execution verification. We perform supervised fine-tuning (SFT) and further reinforcement learning (RL) on both success and failure trajectories with a novel reward design considering code explanation and refinement quality. SFT improves the pass@1 by up to 15.92% and pass@10 by 9.30% over four benchmarks. RL training brings additional up to 3.54% improvement on pass@1 and 2.55% improvement on pass@10. The trained LLMs show iterative refinement ability and can keep refining code continuously. Lastly, our human evaluation shows that the LLMs trained with our framework generate more useful code explanations and help developers better understand bugs in source code.
@@ 1 Introduction
Code generation has become a crucial research task to automatically generate source code based on natural language description [1, 2, 3, 4]. Although the recent Large Language Models (LLMs) have shown impressive capability in code generation, generating the correct code for a complex problem in single attempt is still challenging [5, 6, 7, 8, 9, 10, 11, 12]. This is expected because even for human developers, completing a hard programming problem might need multiple rounds of trial-and-error debugging. Self-debugging capability that allows LLMs to retrospect the incorrect code and make changes to resolve the errors is becoming increasingly important besides their code generation ability.
Existing works [13, 14] investigate off-the-shelf LLMs in the scale of Codex (code-davinci-002[1], GPT-3.5 and GPT-4, and show that these LLMs can self-debug the wrong code they generated via prompting methods in a pipeline of code generation and self-refinement as shown in Figure 1. The user first queries the LLM for a solution for the given programming task and the initial solution from the LLM is verified by executing them against the given unit tests. If the solution passes all the unit tests, it is considered correct. Otherwise, the user collects the unit test feedback and forms a new query to ask the LLM for a refinement. Such a process can iterate until the LLM generates a correct solution or reaches the maximum number of iterations. There are different prompt designs when asking for refinement [13]. Compared with directly asking for a refined solution (referred to as “Code Refinement” in the feedback block), asking LLMs to provide an explanation of the wrong solution and then refine it in a chain-of-thought manner (referred to as “Code Explanation and Refinement” in the feedback block) helps it to better understand the unit test feedback and increases the success rate of providing refined solutions (details in Appendix A.1).
@@
In this work, we propose LeDex, an automated pipeline to collect a high-quality dataset for code explanation and refinement by generating explanation and refinement trajectories, followed by filtering through execution verification. LeDex then leaverages the collected data, using supervised fine-tuning (SFT) to significantly improve LLMs’ ability to explain and refine incorrect code. Additionally, LeDex applies reinforcement learning (RL) with a novel reward design that accounts for explanation semantics and unit test success, leading to better code explanations and corrections. In summary, this work contributes the following:
  • We introduce LeDex, a scalable framework comprising automated data collection, data validation, supervised fine-tuning, and reinforcement learning with novel reward mechanisms to enhance LLMs’ self-debugging capabilities, resulting in more accurate code refinements and insightful code explanations.
  • We experiment LeDex on three backbones (StarCoder-15B, CodeLlama-7B, and CodeLlama-13B) using code refinements and explanations, initially collected from GPT-3.5-Turbo. Supervised fine-tuning notably boosts the models’ ability to diagnose and correct faulty code, achieving up to a 15.92% improvement in pass@1 and a 9.30% increase in pass@10 across four benchmarks.
  • LeDex’s reinforcement learning on top of SFT, uses a novel reward function that incorporates unit test outcomes and semantic analysis of incorrect code explanations. This further enhances performance, with improvements of up to 3.54% in pass@1 and 2.55% in pass@10.
  • LeDex is model-agnostic; notably, CodeLlama-7B trained on data gathered from CodeLlama-34B or even itself achieves up to 8.25% and 2.14% gains in pass@1 and pass@10, demonstrating the generalizability of the approach without reliance on GPT-3.5-Turbo.
@@ 2 Approach
Figure 2 shows the overview of LeDex, including the collection of high-quality code explanation and refinement data, and the training methods. LeDex first collects a code explanation and refinement dataset by querying from pre-trained or instruct models and verifying its responses with execution feedback to filter and obtain high-quality explanation and refinement data (steps 1 and 2 in Figure 2, Section 2.1). Then the high-quality dataset is used for supervised fine-tuning (step 3 in Figure 2, Section 2.2), which significantly improves the model’s performance in explaining the bug and refining the code. Reinforcement learning with execution feedback is used to further guide the model to generate higher quality responses and boost the model performance (step 4 in Figure 2, Section 2.3).
Figure 2: Overview of LeDex.
@@ 2.1 Data collection and verification
Table 1 shows the number of correct (passing all the unit tests) and wrong (failing any unit test) solutions sampled for each dataset. For each wrong solution, we need an explanation of the wrong code and a correct refinement to build the code explanation and refinement dataset. We prompt pre-trained or instruction-LLMs with the problem description, wrong solution, and execution feedback (either error message or failed test case) to ask for an explanation and refinement. We experimented with GPT-3.5-Turbo, CodeLlama-34B, and CodeLlama-7B for data collection. We take GPT-3.5-Turbo as the example in this section, and an example with it is shown in Appendix A.2. We study the generalization of this data collection with different LLMs in Section 4.3.
@@ 4 Result
To evaluate the effectiveness of LeDex, we study the following research questions (RQs) regarding code generation and code refinement capability, iterative refinement ability, approach generalizability, and the quality of the generated code explanations.
@@ 4.1 RQ1: Code generation and refinement capability
We evaluate the models trained with LeDex for their code explanation and refinement ability using four benchmarks: MBPP [3], HumanEval [1], MBPP+ [26], and HumanEval+ [26]. We use pass@k [1] and success refinement rate as the evaluation metric. For the generation of the initial solutions, the models sample 100 solutions per task in the benchmarks (temperature set to 0.8), which are run against the provided test cases. For every incorrect solution that fails any test case, we let the models sample one refinement (and one explanation).
@@ 4.1.1 Pass@k
Table 2 presents the pass@k results across four benchmarks. Overall, fine-tuning the LLMs with our curated dataset of code explanations and refinements leads to substantial improvements in both pass@1 and pass@10 for all three model architectures. For StarCoder-15B and CodeLlama-13B, RL achieves the highest pass@1 and pass@10 scores (bolded) across all four benchmarks. For CodeLlama-7B, RL achieves the best performance in seven out of eight cases, with SFT yielding the highest pass@1 score on the MBPP benchmark.
Table 2: Pass@k of initial and refined solutions on four benchmarks. Each backbone’s best performance on every benchmark is bolded.
@@
Table 3: Overall pass@k on MBPP & HumanEval and MBPP+ & HumanEval+. Blue or red numbers show the improvement or deterioration: SFT is compared to prompting, and RL is compared to SFT.
StarCoder-15B MBPP & HumanEval MBPP+ & HumanEval+
LeDex SFT LeDex RL LeDex SFT LeDex RL
Refine Expl. + Refine Refine Expl. + Refine Refine Expl. + Refine Refine Expl. + Refine
pass@1 54.35 +15.34 54.71 +15.92 56.78 +2.43 56.58 +1.87 49.31 +12.80 49.64 +14.13 50.87 +1.56 51.02 +1.38
pass@10 76.23 +6.58 76.81 +6.30 78.78 +2.55 78.66 +1.58 71.63 +7.27 72.36 +9.30 72.67 +1.04 72.87 +0.51
CodeLlama-7B MBPP & HumanEval MBPP+ & HumanEval+
LeDex SFT LeDex RL LeDex SFT LeDex RL
Refine Expl. + Refine Refine Expl. + Refine Refine Expl. + Refine Refine Expl. + Refine
pass@1 57.21 +14.41 56.75 +14.14 56.62 -0.60 57.41 +0.66 49.59 +11.15 49.95 +11.55 50.73 +1.14 51.74 +1.79
pass@10 78.17 +5.69 78.98 +6.18 79.21 +1.04 80.25 +1.27 72.42 +6.27 71.81 +4.74 74.03 +1.61 74.42 +2.60
CodeLlama-13B MBPP & HumanEval MBPP+ & HumanEval+
LeDex SFT LeDex RL LeDex SFT LeDex RL
Refine Expl. + Refine Refine Expl. + Refine Refine Expl. + Refine Refine Expl. + Refine
pass@1 60.95 +12.23 60.64 +11.68 61.91 +0.96 61.64 +1.00 53.86 +11.14 53.26 +10.51 57.40 +3.54 56.60 +3.34
pass@10 80.83 +3.69 81.29 +3.72 81.09 +0.26 81.74 +0.45 75.21 +4.04 74.87 +3.02 76.50 +1.29 76.00 +1.13
For easier comparison, Table 3 summarizes the overall pass@k results on MBPP and HumanEval, along with the improvements achieved through SFT and RL. The improvements from SFT are compared to direct prompting, while the improvements from RL are relative to SFT.
On MBPP and HumanEval overall, SFT boosts StarCoder-15B’s pass@1 by 15.34% and pass@10 by 6.58% when directly generating code refinements. When incorporating code explanations in a chain-of-thought format, SFT further enhances StarCoder-15B’s performance by 15.92% on pass@1 and 6.30% on pass@10. RL brings an additional 2.43% improvement in pass@1 and 2.55% in pass@10 for direct refinements, and a further 1.87% pass@1 and 1.85% pass@10 increase when generating both code explanations and refinements. Comparable improvements from SFT and RL are observed across the CodeLlama-7B and CodeLlama-13B models as well.
On the MBPP+ and HumanEval+ benchmarks, which feature more rigorous test cases, respectively [26], we observe even greater improvements from RL training on the CodeLlama models. CodeLlama-7B achieves a 1.79% increase in pass@1 and a 2.60% increase in pass@10 for refined solutions with code explanations. CodeLlama-13B shows a 3.54% improvement in pass@1 and a 1.29% improvement in pass@10 for directly generated refinements. These results demonstrate that RL training enables LLMs to produce or refine solutions that are more robust and capable of passing stricter test cases. Additional experiments and detailed case studies can be found in Appendix A.4.1,  A.5.1,  A.5.2, and  A.5.3.
@@ 4.1.2 Success refinement rate
Table 4 presents the refinement success rate for each model backbone across various approaches, averaged over four benchmarks. For StarCoder-15B, the baseline prompting method struggles, achieving only a 6.41% to 6.90% success rate in refining incorrect initial solutions. However, after applying SFT with the high-quality dataset containing code explanations and refinement trajectories, StarCoder-15B demonstrates a notable improvement, raising its refinement success to 16.27% to 16.56%. This increase represents a significant gain of 9.37% to 10.15% over the prompting baseline, showcasing the effectiveness of SFT in enhancing code refinement capabilities by leveraging targeted training data. With further RL, the refinement success for StarCoder-15B improves even more, adding an additional 1.03% to 1.23% over the results from SFT. This final boost highlights the complementary strengths of RL, particularly its capacity to fine-tune model behavior beyond what supervised methods can achieve.
The improvement on CodeLlama-7B and CodeLlama-13B backbones is consistent with that on StarCoder-15B, where RL training eventually achieves the highest success refinement rate with a considerable boost of 1.81 – 3.62%.
Table 4: Success refinement rates over four benchmarks. Blue numbers show the improvement.
@@
Prompt. LeDex SFT LeDex RL Prompt. LeDex SFT LeDex RL
StarCoder-15B 6.90 16.27 +9.37 17.50 +1.23 6.41 16.56 +10.15 17.59 +1.03
CodeLlama-7B 8.65 18.14 +9.49 19.95 +1.81 8.10 17.60 +9.50 20.84 +3.24
CodeLlama-13B 11.64 18.96 +7.32 22.58 +3.62 11.97 20.06 +8.09 23.50 +3.44
@@ 4.2 RQ2: Iterative refinement ability
LLMs have the ability to iteratively self-debug until they arrive at correct solutions. Figure 4 illustrates the overall pass@k of CodeLlama-7B across four benchmarks after up to three rounds of refinements. To simplify the figure, we plot the higher pass@k from either the "Refine" or "Expl. + Refine" approach at each refinement round for each model. Additional results on iterative refinement are provided in Appendix A.4.2.
@@
Both SFT and RL consistently outperform prompting across all three refinement rounds. Even after three rounds, the prompting approach fails to match the pass@k achieved by SFT after just the first round (e.g., 47.53% vs. 56.75% in Figure 4 (a)). These results demonstrate that LLMs trained with our pipeline possess strong iterative refinement capabilities, enabling them to achieve progressively higher pass@k with each additional round of refinement.
Table 5: Pass@k of CodeLlama-7B trained with CodeLlama-34B’s data.
@@
Table 6: Overall pass@k on MBPP & HumanEval and MBPP+ & HumanEval+, trained with CodeLlama-34B’s data. Blue numbers show the improvement.
CodeLlama-7B MBPP & HumanEval MBPP+ & HumanEval+
LeDex SFT LeDex RL LeDex SFT LeDex RL
Refine Expl. + Refine Refine Expl. + Refine Refine Expl. + Refine Refine Expl. + Refine
pass@1 50.01 +7.21 53.15 +10.54 53.20 +0.05 53.64 +0.49 45.62 +7.18 46.03 +7.63 50.76 +5.14 51.31 +5.28
pass@10 76.88 +4.40 77.00 +4.20 77.42 +0.42 77.33 +0.33 69.57 +3.42 69.74 +2.67 71.65 +2.08 71.98 +2.24
4.3 RQ3: Generalizability of approach
4.3.1 Data collection using open source LLM
To demonstrate the generalizability of LeDex, particularly the independence of our data collection process from GPT-3.5-Turbo, we substitute GPT-3.5-Turbo with CodeLlama-34B for data collection. As CodeLlama-34B is a pre-trained model, we incorporate few-shot examples in the prompts to guide the generation of incorrect code explanations and refinements. All other processes remain unchanged.
Table 5 presents the pass@k results for CodeLlama-7B trained on data collected from CodeLlama-34B, with Table 6 providing an overall comparison. Although SFT achieves slightly smaller improvements (around 1–3% lower than with GPT-3.5-Turbo data), it still yields notable gains in overall pass@1 and pass@10. Additionally, we observe that RL training further enhances performance on the MBPP+ and HumanEval+ benchmarks, with pass@1 improving by 5.28% and pass@10 by 2.24%. These results demonstrate the generalizability of LeDex and further suggest that collecting data from a more powerful LLM can lead to better training outcomes within our framework. Additional results can be found in Appendix A.4.3.
Table 7: Pass@k of CodeLlama-7B trained with self-bootstrapped data.
Models Approaches MBPP Humanval MBPP+ HumanEval+
pass@1 pass@10 pass@1 pass@10 pass@1 pass@10 pass@1 pass@10
CodeLlama-7B Init. 45.83 69.24 39.85 68.83 41.78 61.77 33.25 61.50
LeDex SFT Refine 52.37 74.63 47.04 74.58 46.26 66.39 40.15 67.15
Expl. + Refine 51.80 74.99 45.70 74.72 45.94 65.77 39.10 67.33
Init. 46.28 68.87 39.90 69.49 41.61 61.29 33.66 62.17
LeDex RL Refine 52.84 74.46 48.37 75.54 46.28 65.86 41.54 68.14
Expl. + Refine 52.34 74.60 46.90 75.70 46.10 65.99 40.79 68.50
Table 8: Overall pass@k on MBPP & HumanEval and MBPP+ & HumanEval+, trained with self-bootstrapped data. Blue or red numbers show the improvement or deterioration.
CodeLlama-7B MBPP & HumanEval MBPP+ & HumanEval+
LeDex SFT LeDex RL LeDex SFT LeDex RL
Refine Expl. + Refine Refine Expl. + Refine Refine Expl. + Refine Refine Expl. + Refine
pass@1 51.05 +8.25 50.29 +7.68 51.74 +0.69 51.00 +0.71 43.77 +5.33 43.16 +4.76 44.35 +0.58 43.94 +0.78
pass@10 74.62 +2.14 74.92 +2.12 74.73 +0.11 74.87 -0.05 66.70 +0.55 66.40 -0.67 66.79 +0.09 67.01 +0.61